# A Systematic Survey on Large Language Models for Algorithm Design

### v4

@@ Abstract.
Algorithm design is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and promising solutions. In just a few years, this integration has yielded remarkable progress in areas ranging from combinatorial optimization to scientific discovery. Despite this rapid expansion, a holistic understanding of the field is hindered by the lack of a systematic review, as existing surveys either remain limited to narrow sub-fields or with different objectives. This paper seeks to provide a systematic review of algorithm design with LLMs. We introduce a taxonomy that categorises the roles of LLMs as optimizers, predictors, extractors and designers, analyzing the progress, advantages, and limitations within each category. We further synthesize literature across the three phases of the algorithm design pipeline and across diverse algorithmic applications that define the current landscape. Finally, we outline key open challenges and opportunities to guide future research.
@@ 1. Introduction
Year Survey LLM Algorithmic Design Stages Algorithm Type
@@
Figure 1. Overview of the Survey Structure: Methodology, Taxonomy of LLM Roles, Stages of LLM‑assisted Algorithm Design, Applications, Open Challenges and Future Directions.
@@ 3.3. LLM as Extractor (LLMaE)
Beyond embedding-based feature extraction, LLMs excel in text comprehension and knowledge extraction, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional feature extraction methods. For example, wu2024large utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. du2024mixture propose a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem. Additionally, memduhouglu2024enriching use LLM to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which have previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
@@ 3.4. LLM as Designer (LLMaD)
EoH (liu2024evolution; liu2023algorithm) presents an early attempt to adopt the LLM as a designer for Automated Heuristic Design (AHD). It uses both heuristic ideas and code implementations to represent heuristics and adopts LLM in an evolutionary framework to create, combine, and revise the heuristics. While it is original proposed for heuristic design, it has been applied on different algorithm design tasks including combinatorial optimization problems (liu2024evolution; yao2024multi; liu2025eoh), Bayesian optimization (yao2024evolve), image adversary attack (guo2024autoda), and edge server task scheduling (wang2024ts), among others. Moreover, LLaMEA (van2024llamea; van2024loop) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (ye2024reevo) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. In addition to evolutionary search framework, recent attempts also adopt other frameworks such as large neighborhood search (xie2025llm) and Monte Carlo Tree Search (MCTS) (zhengmonte) for effective explore the algorithm space. Unlike previous studies that focus on optimizing a single performance criterion, MEoH (yao2024multi) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space.
@@ 4.2. Stage II: Algorithm Implementation
A common pattern for algorithmic code optimization is closed-loop refinement, where execution feedback informs subsequent code edits and integration choices (zhang2024understanding). Evolutionary search is the most prevalent framework for this purpose, maintaining a population of algorithms that are progressively refined by an LLM (zhang2024understanding). Variants of this simple evolutionary approach have emerged. For instance, ReEvo combines short- and long-term reflection with information on history evaluated results to guide algorithm optimization (ye2024reevo), while HSEvo adopts diversity control strategies with two population diversity measurements to enhance the optimization process (yao2025evolution). Recently, non-evolutionary search frameworks, such as large neighborhood search and Monte Carlo tree search (zhengmonte), have also been demonstrated to be effective. Moreover, instead of prompting LLMs to optimize algorithm implementations, recent attempts have explored fine-tuning LLMs, in both offline (liu2025fine) and online manners (huang2025calm), to learn a preference for generating better-performing algorithms during algorithm optimization for specific design tasks.
@@ 4.3. Stage III: Algorithm Evaluation
In addition to generation, LLMs are also being employed to analyze algorithmic outcomes and explain their behavior. d2024exploring empirically analyzed and compared the quality of explanations provided by three different LLMs for seven state-of-the-art quantum algorithms. Their findings indicated that while the explanations were consistent across multiple iterations, there remains a significant gap for improvement in quality, and the results were highly sensitive to the specific prompts used. In a different application, chacon2024large integrated LLMs into STNWeb, an algorithm analysis tool. The LLM generates extensive written reports, complemented by automatically generated plots, which enhances the user experience and lowers the barrier to adoption for the broader research community.
@@ 5.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), automated algorithm heuristics design has been a significant area of interest for a long time. The Traveling Salesman Problem (TSP) stands out as one of the most renowned CO problems, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (liu2023algorithm), ReEvo (ye2024reevo) and EoH (liu2024evolution). Differently, OPRO (yang2024large) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (liu2024largelmea) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (huang2024multimodal) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance. In addition to routing problems, other combinatorial optimization problems that have also been investigated include cap set (romera2024mathematical), bin packing (liu2024evolution), flow shop scheduling (liu2024evolution), hybrid job shop scheduling (li2025llm) and social networks problems (sartori2024metaheuristics).
@@ 5.1.2. Continuous Optimization
In various real-world applications, the cost of evaluating the objective functions can be very expensive, which greatly limits the evaluation budget in the optimization process (frazier2016bayesian). Bayesian optimization (BO) stands out as a sample-efficient method, it typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to carefully select potential solutions. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (zhang2023using) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. Furthermore, LLAMBO (liu2024largellambo) incorporates LLM capabilities to enhance BO efficiency, in which three specific enhancements throughout the BO pipeline have been systematically investigated on tasks selected from HPOBench (eggensperger2021hpobench). Instead of utilizing LLMs for direct solution generation, BO-LIFT* (ramos2023bayesian) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (dinh2022lift) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (yao2024evolve) introduces a novel paradigm integrating LLMs within EC framework to design AFs automatically for cost-aware BO, the approach showcases remarkable efficiency and discovers novel ideas not previously explored in existing literature on AF design. Similarly, FunBO (aglietti2025funbo) found novel and well-performing AFs for BO by extending FunSearch (romera2024mathematical), the discovered AFs are evaluated on various synthetic and HPO benchmarks in and out of the training distribution.
@@ 5.3. Scientific and Symbolic Domains
This subsection is dedicated to exploring LLM-based scientific discoveries related to algorithm designs. Table. 4 lists the related works in this domain.
@@ 5.3.3. Mechanics
MechAgents (ni2024mechagents) introduces a class of physics-inspired generative machine learning platforms that utilize multiple LLMs to solve mechanics problems, such as elasticity, by autonomously collaborating to write, execute, and self-correct code using finite element methods. Moreover, du2024large adopt LLMs in automatically discovering governing equations from data, utilizing the models’ generation and reasoning capabilities to iteratively refine candidate equations. This approach, tested on various nonlinear systems including the Burgers, Chafee–Infante, and Navier–Stokes equations, demonstrates a superior ability to uncover correct physical laws and shows better generalization on unseen data compared to other models. AutoTurb (zhang2024autoturb), on the other hand, adopts LLMs in an evolutionary framework to automatically design and search for turbulence model in computational fluid dynamics. Furthermore, buehler2023melm study LLM-based methods for forward and inverse mechanics problems including bio-inspired hierarchical honeycomb design, carbon nanotube mechanics, and protein unfolding.
@@ 6.5. Benchmarking
While related benchmarks have been developed for mathematical reasoning (lin2024atg), planning (valmeekam2024planbench), and classical algorithmic tasks (velivckovic2022clrs), they do not fully capture the nuances of LLM-driven design. Recently, efforts have begun to develop benchmarks specifically for LLM-driven algorithm design. For example, sun2025co and chen2025heurigym introduce benchmark suites focused on combinatorial optimization. van2025blade developed a benchmark for designing heuristics for black-box optimization. Broadening the scope, liu2024llm4ad proposed a collection of tasks spanning diverse domains, from optimization and machine learning to scientific discovery.
@@ 7. Conclusion
8. Acknowledgments
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China [GRF Project No. CityU11217325].

### v5

@@ Abstract.
Algorithm design is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and promising solutions. In just a few years, this integration has yielded remarkable progress in areas ranging from combinatorial optimization to scientific discovery. Despite this rapid expansion, a holistic understanding of the field is hindered by the lack of a systematic review, as existing surveys either remain limited to narrow sub-fields or with different objectives. This paper seeks to provide a systematic review of algorithm design with LLMs. We introduce a taxonomy that categorises the roles of LLMs as optimizers, predictors, extractors and designers, analyzing the progress, advantages, and limitations within each category. We further synthesize literature across the three phases of the algorithm design pipeline and across diverse algorithmic applications that define the current landscape. Finally, we outline key open challenges and opportunities to guide future research. To support future research and collaboration, we provide an accompanying repository at: https://github.com/FeiLiu36/LLM4AlgorithmDesign.
@@ 1. Introduction
Year Survey LLM Algorithmic Design Stages Algorithm Types
@@
Figure 1. Overview of the Survey Structure: Scope, Taxonomy of LLM Roles, Stages of Algorithm Design, Applications, Open Challenges and Future Directions.
@@ 3.3. LLM as Extractor (LLMaE)
Beyond embedding-based feature extraction, LLMs excel in text comprehension and knowledge extraction, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional feature extraction methods. For example, wu2024large utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. du2024mixture propose a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem. Additionally, memduhouglu2024enriching use LLM to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which have previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts. Beyond extracting problem features, LLMs are employed to mine relevant knowledge to inform and enhance algorithm design. In HiFo-Prompt (chen2025hifo), for example, knowledge is distilled into reusable design principles that guide the design process.
@@ 3.4. LLM as Designer (LLMaD)
EoH (liu2024evolution; liu2023algorithm) presents an early attempt to adopt the LLM as a designer for Automated Heuristic Design (AHD). It uses both heuristic ideas and code implementations to represent heuristics and adopts LLM in an evolutionary framework to create, combine, and revise the heuristics. While it is original proposed for heuristic design, it has been applied on different algorithm design tasks including combinatorial optimization problems (liu2024evolution; yao2024multi; liu2025eoh), Bayesian optimization (yao2024evolve), image adversary attack (guo2024autoda), and edge server task scheduling (wang2024ts), among others. Moreover, LLaMEA (van2024llamea; van2024loop) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (ye2024reevo) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. HSEvo (dat2025hsevo) and PartEvo (hu2025partition) integrate different diversity control strategies in evolutionary search framework to enhance the search. In addition to evolutionary search framework, recent attempts also adopt other frameworks such as large neighborhood search (xie2025llm) and Monte Carlo Tree Search (MCTS) (zhengmonte) for effective explore the algorithm space. Unlike previous studies that focus on optimizing a single performance criterion, MEoH (yao2024multi) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space.
@@ 4.2. Stage II: Algorithm Implementation
A common pattern for algorithmic code optimization is closed-loop refinement, where execution feedback informs subsequent code edits and integration choices (zhang2024understanding). Evolutionary search is the most prevalent framework for this purpose, maintaining a population of algorithms that are progressively refined by an LLM (zhang2024understanding). Variants of this simple evolutionary approach have emerged. For instance, ReEvo (ye2024reevo) combines short- and long-term reflection with information on history evaluated results to guide algorithm optimization, while HSEvo (dat2025hsevo) adopts diversity control strategies with two population diversity measurements to enhance the optimization process. Recently, non-evolutionary search frameworks, such as large neighborhood search and Monte Carlo tree search (zhengmonte), have also been demonstrated to be effective. Moreover, instead of prompting LLMs to optimize algorithm implementations, recent attempts have explored fine-tuning LLMs, in both offline (liu2025fine) and online manners (huang2025calm), to learn a preference for generating better-performing algorithms during algorithm optimization for specific design tasks.
@@ 4.3. Stage III: Algorithm Evaluation
In addition to generation, LLMs are also being employed to analyze algorithmic outcomes and explain their behavior. d2024exploring empirically analyze and compare the quality of explanations provided by three different LLMs for seven state-of-the-art quantum algorithms. Their findings indicate that while the explanations were consistent across multiple iterations, there remains a significant gap for improvement in quality, and the results were highly sensitive to the specific prompts used. In a different application, chacon2024large integrate LLMs into STNWeb, an algorithm analysis tool. The LLM generates extensive written reports, complemented by automatically generated plots, which enhances the user experience and lowers the barrier to adoption for the broader research community.
@@ 5.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), automated algorithm heuristics design has been a significant area of interest for a long time. The Traveling Salesman Problem (TSP) stands out as one of the most renowned CO problems, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Some recent work leverages LLMs to evolve algorithms within evolutionary computation framework, such as EoH (liu2024evolution) and ReEvo (ye2024reevo). Differently, OPRO (yang2024large) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (liu2024largelmea) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (huang2024multimodal) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance. In addition to routing problems, other combinatorial optimization problems that have also been investigated include cap set (romera2024mathematical), bin packing (liu2024evolution), flow shop scheduling (liu2024evolution), hybrid job shop scheduling (li2025llm) and social networks problems (sartori2024metaheuristics).
@@ 5.1.2. Continuous Optimization
In various real-world applications, the cost of evaluating the objective functions can be very expensive, which greatly limits the evaluation budget in the optimization process (frazier2016bayesian). Bayesian optimization (BO) stands out as a sample-efficient method, it typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to carefully select potential solutions. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (zhang2023using) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. Furthermore, LLAMBO (liu2024largellambo) incorporates LLM capabilities to enhance BO efficiency, in which three specific enhancements throughout the BO pipeline have been systematically investigated on tasks selected from HPOBench (eggensperger2021hpobench). Instead of utilizing LLMs for direct solution generation, BO-LIFT* (ramos2023bayesian) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (dinh2022lift) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (yao2024evolve) introduces a novel paradigm to design AFs automatically for cost-aware BO. The approach showcases remarkable efficiency and discovers novel ideas not previously explored in existing literature on AF design. Similarly, FunBO (aglietti2025funbo) found novel and well-performing AFs for BO by extending FunSearch (romera2024mathematical). The discovered AFs are evaluated on various synthetic and HPO benchmarks in and out of the training distribution.
@@ 5.3. Scientific and Symbolic Domains
This subsection is dedicated to exploring LLM-based scientific discoveries. Rather than focusing on classic algorithm design, the works discussed here demonstrate how LLMs function as algorithmic engines to create search strategies, predict performance, and generate symbolic expressions to solve complex scientific problems. Table. 4 lists the related works in this domain.
@@ 5.3.3. Mechanics
MechAgents (ni2024mechagents) introduces a class of physics-inspired generative machine learning platforms that utilize multiple LLMs to solve mechanics problems, such as elasticity, by autonomously collaborating to write, execute, and self-correct code using finite element methods. Moreover, du2024large adopt LLMs in automatically discovering governing equations from data, utilizing the models’ generation and reasoning capabilities to iteratively refine candidate equations. This approach, tested on various nonlinear systems including the Burgers, Chafee–Infante, and Navier–Stokes equations, demonstrates a superior ability to uncover correct physical laws and shows better generalization on unseen data compared to other models. AutoTurb (zhang2024autoturb), on the other hand, adopts LLMs in an evolutionary framework to automatically design and search for turbulence model in computational fluid dynamics. Furthermore, buehler2023melm study LLM-based methods for forward and inverse mechanics problems, including bio-inspired hierarchical honeycomb design, carbon nanotube mechanics, and protein unfolding.
@@ 6.5. Benchmarking
While related benchmarks have been developed for mathematical reasoning (lin2024atg), planning (valmeekam2024planbench), and classical algorithmic tasks (velivckovic2022clrs), they do not fully capture the nuances of LLM-driven design. Recently, efforts have begun to develop benchmarks specifically for LLM-driven algorithm design. For example, sun2025co and chen2025heurigym introduce benchmark suites focused on combinatorial optimization. van2025blade developed a benchmark for designing metaheuristics for black-box optimization. Broadening the scope, liu2024llm4ad proposed a collection of tasks spanning diverse domains, from optimization and machine learning to scientific discovery.